Please answer this question:
You have a pool that measures 8 yards by 10 yards. You want to build a deck around all four siddes of the pool and you want to increase the total area to 120 square yards. The deck is the same width on both sides forming a perfect rectangle. What is the width of the deck?

Answer by checkley77(12844)

8*10=80 yds^62 is the area of the pool.
(8+2x)(10+2x)=120
80+20x+16x+4x^2=120
4x^2+36x+80-120=0
4x^2+36x-40=0
4(x^2+9x-10)=0
4(x+10)(x-1)=0
x-1=0
x=1 yd. is the width of the deck.
Proof:
(8+2*1)(10+2*1)=120
(8+2)(10+2)=120
10*12=120
120=120